OpenStudy (anonymous):

which polynomials are perfect square trinomials: 9x^2 − 6x + 1...........36b^2 − 24b + 8..........16x^2 + 24x+ 9...........4a^2 − 10a + 25

OpenStudy (campbell_st):

well a perfect trinomial... has 3 terms... (ax + b)^2 = a^2x^2 + 2abx + b^2 or (ax - b)^2 = a^2b^2 -2abx + b^2 so look for 1st and last terms that are squares and a middle term that is double the product of the 1st and last. so looking at the 1st expression 9x^2 = (3x)^2 1 = (1)^2 so double the product will be -6x or 6x which you seem to have so the 1st expression can be written as (3x - 1)^2 = 9x^2 - 6x + 1 hope it makes sense and helps

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

the 2nd equation isn't a perfect square as 8 isn't a square number